# Topological characterization of magnon-polaron bands and thermal Hall conductivity in a frustrated kagome antiferromagnet

### v1

@@ Abstract
Spin-phonon coupling and its efficacy in inducing multiple topological phase transitions in a frustrated kagome antiferromagnet have not been addressed in the literature. To this end, we study the ramifications of invoking optical phonons in a non-collinear antiferromagnet via two different kinds of coupling mechanisms, namely, a local and a non-local one, which are distinct in their macroscopic origin. In the case of a local spin-phonon coupling, a single phonon mode affects the magnetic interactions, whereas in the non-local case, two neighbouring phonon modes are involved in the energy renormalization, and it would be worthwhile to compare and contrast the two. To tackle these phonons, we propose an analytical approach through a canonical spin-Peierls transformation applied to magnons, that renders a hybridization between the magnons and the phonon modes and yields a magnon-polaron quasiparticle state. In both the coupling regimes, the robust support of the topological signatures is derived systematically from the bulk and edge spectral properties of the magnon-polaron bands that are characterized by their corresponding Chern numbers. Thereafter, we witness transitions from one topological phase to another solely via tuning the spin-phonon coupling strength. Moreover, these transitions significantly impact the behavior of the thermal Hall conductivity, which aid in discerning distinct topological phases. Additionally, the explicit dependencies on the temperature and the external magnetic field are explored in inducing topological phase transitions associated with the magnon-polaron bands. Thus, our work serves as an ideal platform to probe the interplay of frustrated magnetism and polaronic physics.
@@ I INTRODUCTION
Of late, geometrically frustrated kagome antiferromagnets (KAFM) have been proven to be an important candidate to host exotic QSL features [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. Characterized by various elementary excitations, the theoretical studies have speculated on ground state QSL markers of the kagome Heisenberg antiferromagnet, such as the $\mathbb{Z}_{2}$ [45, 36, 37, 41, 46], Dirac [39, 42, 43, 44], topologically entangled [38], and chiral [47, 24] spin liquid states, and more recently, spinon pair-density wave state [48]. Although, among the QSLs chiral spin liquids break the time-reversal symmetry (TRS) and result in formation of Chern bands, it is unclear whether the QSL ground state of the KAFM is gapped [35, 36, 37, 41, 37, 39, 43]. In addition to these fundamental issues, a KAFM harbors worthwhile quantum signatures induced by DMI [49, 50, 51, 52, 53, 54, 20, 21, 22, 23, 24]. The DMI can intrinsically induce long-range order (LRO), commonly termed $\bm{Q}=0$ magnetic order in frustrated kagome magnets, which is otherwise inaccessible at low temperatures. Hence, up to a critical order, DMI suppresses the QSL phase of the KAFM [51]. Past studies have evidenced the LRO in coplanar or noncollinear frustrated KAFM at specific regimes of temperature [55, 56, 49, 57, 50, 51]. However, theoretical predictions on noncoplanar KAFM structures, such as pyrochlore thin film [58], jarosites [59, 60, 61, 56, 49, 50], namely $\text{KFe}_{3}\text{(OH)}_{6}(\text{SO}_{4})_{2}$, $\text{AgFe}_{3}\text{(OH)}_{6}(\text{SO}_{4})_{2}$, etc., Vesignieite $\text{BaCu}_{3}\text{V}_{2}\text{O}_{8}\text{(OH)}_{2}$ [53], herbertsmithite $\text{ZnCu}_{3}\text{(OH)}_{6}\text{Cl}_{2}$ [62], and noncollinear canted configurations induced by a magnetic field [63, 64, 65] have been probed to discern MTHE experimentally. Due to the inherent noncoplanarity in several of these KAFMs, a DMI-induced scalar spin chirality, $\chi=\bm{S}_{i}.(\bm{S}_{j}\times\bm{S}_{k})$ originates which instills a nontrivial magnon topology [49, 50, 51, 54, 20, 21] and MTHE [20, 21, 22, 23, 24]. While a field-induced noncoplanar spin chirality is solely capable of manifesting MTHE in KAFMs, with no mandatory requirement of DMI [20]. In contrast, with no magnetic field, the in-plane DMI, has been reported to stabilize a canted KAFM configuration with a finite spin chirality that significantly modifies the magnon thermal Hall conductivity [21] corresponding to each of the Chern (magnon) bands. Moreover, KAFMs host a “spinon quantum spin Hall effect” in the time-reversal symmetric gapped QSL phase and facilitate a continuous phase transition into this phase with the inclusion of an extended nearest neighbour hopping that melts the DMI-induced $\bm{Q}=0$ ordering [24]. Hence, this gapped QSL phase with oppositely moving spinon chiral edges (possessing spin Chern numbers, $C_{\uparrow,\downarrow}=\pm 1$) renders itself as a magnetic analogue of the $\mathbb{Z}_{2}$ topological insulator. However, even with all of these proposals being accounted for, there is still enough scope to explore quasiparticle excitonic topology and MTHE of the KAFM magnon bands.
@@
Intriguingly, other variants of quasiparticles, such as phonons (associated with the lattice vibrations), also corroborate nontrivial topological phenomena. Akin to phonon-induced band topology in electronic systems [66, 67, 68, 69, 70, 71, 72, 73], these vibrational modes can hybridize with magnons, resulting in magnon-phonon (also termed magnon-polaron) quasiparticle coupled topological signatures in magnetically ordered systems [74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93]. To mention a few, the hybridization effects of the magnon and phonon bands are mostly explored in square [77, 79, 82], honeycomb [81, 86, 87, 89, 90, 91] lattices, and two-dimensional (2D) antiferromagnets [75, 76, 79, 82, 86, 88, 90]. It has been claimed that hybridization of these quasiparticles can occur with (via breaking the mirror symmetry) [80, 77, 81, 86, 89] or without [78, 76, 74] DMI or through dipolar interactions [74], which results in anticrossing of these excitonic bands accompanied by a nontrivial Berry curvature that induces magnon-polaron thermal Hall effect. The hybridization in the absence of any DMI is usually explained by magnetostriction or magnetoelastic wave [94, 78, 76], which forms a magnon-polaron state mainly through the acoustic phonon modes. While the role of optical phonon modes has also been examined in manipulating the Chern numbers of the hybridized magnon-polaron bands [84, 86], such studies are non-existance in the context of a kagome lattice. Moreover, magnon-phonon topological physics in the kagome structure is not well established in the literature, especially on a KAFM configuration. Hence, a closer inspection of the magnon-optical phonon coupled quasiparticle scenario (in terms of magnon-polaron bound state) on a KAFM may unravel important topological properties of frustrated magnetism, which, to the best of our knowledge, has not been reported yet.
@@
In this study, we explore the effects of magnon-polaron excitations on a 2D noncoplanar KAFM (as schematized in Fig. 1(a)) in the presence of an out-of-plane DMI ($D$) and an external magnetic field ($B_{0}$). Additionally, we formulate our magnon model with the nearest neighbour (NN) and the next-nearest neighbour (NNN) Heisenberg exchange terms, namely $J_{1}$ and $J_{2}$, respectively. However, we mention that the in-plane DMI components being negligibly small due to the dominant out-of-plane DMI in most kagome materials [55, 62, 56, 51], we drop the in-plane DMI terms here. To inspect a frustrated magnetism on our system, the noncoplanarity or the canted configuration (away from $120^{\circ}$ Néel ordering) of KAFM necessitates an externally applied out-of-plane magnetic field that solely generates a nontrivial magnon topology and MTHE [20]. In general, the interaction between quasiparticles in such a KAFM model can be intricate and hence requires involved techniques to derive an effective magnon picture. Here, we wish to investigate how the onsite (local) magnon modes couple to the optical phonon modes, resulting in a magnon-polaron state, and examine its impact on the magnon topology and the relative thermal Hall features of the entire KAFM structure. Focusing on this primary goal, we first consider that the optical phonon modes can couple with the spin (magnon) modes through two distinct mechanisms: (i) magnon modes at a particular lattice site interact with the same onsite phonon modes [95], named “local spin-phonon” coupling with a strength $g^{l}$, and (ii) local magnon interaction that involves phonons present at two NN lattice sites [95, 84, 78], dubbed “non-local spin-phonon” coupling with a strength $g^{nl}$. The purpose of dealing with optical (rather than the acoustic) phonon branches is to incorporate a stronger perturbation on the local atomic environment, and hence it ascertains more efficient coupling between the spin excitations and phonons [84], which can be probed via photo-excitation experiments. Furthermore, going beyond the adiabatic (high frequency) regime enables phonons to acquire a finite energy that stabilizes the KAFM configuration even at zero temperature, although a constraint on the phonon frequency should be imposed to avoid instability of the magnon-polaron bands. In addition, as optical modes are largely non-dispersive, it is convenient to model the phonon bands generically at small momenta [96] that are relevant for gap-closing transitions of the KAFM bands. These justify the rationale for taking momentum-independent optical phonons into account. However, analytical methods for tackling such a coupling in a KAFM have not been proposed yet. Therefore, here we provide an analytical approach in the spirit of the spin-Peierls transformation method [95] to yield a finite temperature magnon-polaron hybridized bound state by completely decoupling these two distinct couplings on a KAFM, and then analyze numerically its topological properties and the thermal Hall conductivity. Although hitherto proposals on the kagome ferromagnet [81] and the kagome spin ice [93] have explained the effects of phononic excitations, the salient topology of the magnon-phonon hybridization in a frustrated geometry, such as a KAFM has not been addressed till now. Therefore, the present study may hint at a richer topological scenario and bridge the gap between frustrated magnetism and magnon-polaron quasiparticle physics. Thus far, the following pertinent questions arise: (i) Can a topological transition be induced on the KAFM magnon-polaron bands solely by tuning the spin-phonon coupling strength? (ii) How will the magnon-polaron band topology differ in local and non-local interaction regimes, both qualitatively and quantitatively? (iii) What will the response of the topological thermal Hall conductivity be to such magnon-polaron hybridization in both these regimes?
@@
To address these points, we organize the remainder of the paper as follows. We begin by mathematically formulating the noncoplanar KAFM magnon model, and construct the momentum ($k$) space magnon-polaron Hamiltonian, followed by the exact decoupling of the aforementioned coupling through a canonical transformation in Sec. II. Subsequently, in Sec. III, we describe the numerical results in the following sequence. The bulk spectral properties accompanied by relative transitions in the Chern number are examined in Sec. III.1 via varying both the local and non-local coupling strengths distinctly, while the edge state features of the topological magnon-polaron bands are analyzed in Sec. III.2. Thereafter, the nature of the topological thermal Hall conductivities exhibiting the transitions is elaborately discussed in Sec. III.3. Finally, we summarize and conclude our findings in Sec. IV.
@@
Figure 1: Schematic diagram of a kagome antiferromagnet, (a) where red, blue, and green arrows refer to the spin orientation of A, B, and C sublattices, respectively, and $\mathbf{\delta}_{1}$, $\mathbf{\delta}_{2}$ and $\mathbf{\delta}_{3}$ are three lattice vectors. (b) The direction of the out-of-plane DMI is shown, it acts downward from site 1 to 2, and upward from site 2 to 3. (c) In the presence of an external magnetic field, the spins become canted with an angle $\eta$ with respect to $x$-$y$ plane.
@@ II.1 Magnon Hamiltonian
where $J_{1}(>0)$ and $J_{2}(>0)$ are the NN and the NNN antiferromagnetic exchange interaction strengths, as mentioned earlier. Additionally, $D$ is the coefficient of an out-of-plane DMI with a positive chirality acting between two neighbouring spins, where $\nu_{ij}=\mp 1$ depending on their lattice positions, that is, negative (positive) for an upper (lower) triangle (see Fig. 1(b)) [20]. Finally, the Zeeman term involves the external transverse magnetic field $B_{0}$ that eventually generates the canted orientations of spins and further aids in inducing the nontrivial topology in the system. The lattice vectors connecting two same sublattice sites are denoted by, $\delta_{1}=a_{0}(1,\sqrt{3})$, $\delta_{2}=a_{0}(1,-\sqrt{3})$, $\delta_{3}=a_{0}(2,0)$.
Further, we consider non-collinear spin configurations with antiferromagnetic environment, where, neglecting all the spin fluctuations, the classical spin vectors at three different sublattice sites (A, B, C) (as displayed in Fig. 1(a)) are given by
@@
Here, $\eta$ denotes the canting angle between the spin vectors and the lattice plane, as shown in Fig. 1(c)). In the absence of an external magnetic field, the spin configuration exhibits a $120^{\circ}$ Néel (coplanar) order, with $\eta=0$. However, in the presence of an external magnetic field ($B_{0}\neq 0$), $\eta$ acquires a finite value, which can be determined by minimizing the classical energy ($\partial E_{\text{cl}}/\partial\eta=0$) corresponding to $H_{\text{s}}$ (Eq. (2)). This yields the relation $\sin{\eta}=B_{0}/B_{s}$, with $B_{s}$ being the saturation magnetization given by, $B_{s}=2\sqrt{3}[D+\sqrt{3}(J_{1}+J_{2})]$, corresponding to the ground-state spin configuration as reported in literature [20, 21].
Due to this canted spin orientation, the spin quantization axes at different sublattice sites not only differ from one another but also deviate from the global $z$-axis. Therefore, to establish a relation between the global spin operator $\mathbf{S}_{i}$ and the local spin operator $\tilde{\mathbf{S}}_{i}$, we need to introduce sublattice-dependent rotation matrix, that can be expressed as
@@
where $\tilde{S}_{i}^{+}=\tilde{S}_{x}+i\tilde{S}_{y}$, $\tilde{S}_{i}^{-}=\tilde{S}_{x}-i\tilde{S}_{y}$ and $a_{i}^{\dagger}(a_{i})$ is the creation (annihilation) operator corresponding to the A sublattice sites. Similarly, for the B and C sublattice sites, the creation (annihilation) operators are denoted by $b_{i}^{\dagger}$ ($b_{i}$) and $c_{i}^{\dagger}$ ($c_{i}$), respectively. In our case, these operators follow the same transformation as given in Eq. (5). In the frustrated antiferromagnetic system, the spin orientations at the different sublattice sites are specified in Eq. (4). By applying the spin-boson transformation defined in Eq. (5) to the spin Hamiltonian in Eq. (2), the Hamiltonian $H_{\text{s}}$ can be expressed in the magnon basis, and is given by
$H_{\text{s}}$ $=$ $S\sum_{r}\left[4(J_{1}+J_{2})\mathcal{E}+4D{\mathcal{E}}_{1}+\mu_{0}B_{0}\sin{\eta}\right]$ (6)
$\left[a_{r}^{\dagger}a_{r}+b_{r}^{\dagger}b_{r}+c_{r}^{\dagger}c_{r}\right]+\left[J_{1}\!(P\!+\!iQ)\!-\!D(P_{1}\!+\!iQ_{1})\right]$
$\left[c_{r}^{\dagger}a_{r}+c_{r}^{\dagger}a_{r+\delta_{1}}+b_{r}^{\dagger}c_{r}+b_{r}^{\dagger}c_{r+\delta_{2}}+a_{r}^{\dagger}b_{r}+a_{r+\delta_{3}}^{\dagger}b_{r}\right]$
$+J_{2}(P+iQ)[a_{r}^{\dagger}b_{r-\delta_{2}}+a_{r+\delta_{1}}^{\dagger}b_{r}+b_{r+\delta_{1}}^{\dagger}c_{r}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+b_{r}^{\dagger}c_{r+\delta_{3}}+c_{r+\delta_{2}}^{\dagger}a_{r}+c_{r}^{\dagger}a_{r+\delta_{3}}]$
$+\left[J_{1}R-DR_{1}\right][c_{r}^{\dagger}a_{r}^{\dagger}+c_{r}^{\dagger}a_{r+\delta_{1}}^{\dagger}+b_{r}^{\dagger}c_{r}^{\dagger}+b_{r}^{\dagger}c_{r+\delta_{2}}^{\dagger}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+a_{r}^{\dagger}b_{r}^{\dagger}+a_{r+\delta_{3}}^{\dagger}b_{r}^{\dagger}]$
$+J_{2}R[a_{r}^{\dagger}b_{r-\delta_{2}}^{\dagger}+a_{r+\delta_{1}}^{\dagger}b_{r}^{\dagger}+b_{r+\delta_{1}}^{\dagger}c_{r}^{\dagger}+b_{r}^{\dagger}c_{r+\delta_{3}}^{\dagger}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+c_{r+\delta_{2}}^{\dagger}a_{r}^{\dagger}+c_{r}^{\dagger}a_{r+\delta_{3}}^{\dagger}]+h.c.,$
where $r$ denotes the position of a unit cell and the sum runs over all the unit cells. The operators $a_{r}^{\dagger}$ $(a_{r})$, $b_{r}^{\dagger}$ $(b_{r})$ and $c_{r}^{\dagger}$ $(c_{r})$ are the creation (annihilation) operators corresponding to sublattices A, B, and C, respectively, located at position $r$, corresponding to the triangle highlighted in yellow in Fig. 1(a). Moreover, each unit cell contains six bonds, comprising of one upper and one lower triangle associated with that unit cell. Further, the remaining quantities of Eq. (6) are defined as

### v2

@@ Abstract
Spin-phonon coupling and its efficacy in inducing multiple topological phase transitions in a frustrated kagome antiferromagnet have been rare in literature. To this end, we study the ramifications of invoking optical phonons in such a system via two different coupling mechanisms, namely, a local and a non-local one, which are distinct in their microscopic origin. In case of the local spin-phonon coupling, a single phonon mode affects the magnetic interactions, whereas in the non-local case, two neighbouring phonon modes are involved in the energy renormalization, and it would be worthwhile to compare and contrast between the two. To tackle these phonons, we propose an analytic approach involving a canonical spin-Peierls transformation applied to magnons. The formalism renders a hybridization between the magnons and the phonon modes, yielding magnon-polaron quasiparticles. In both the coupling regimes, validations for the topological signatures are systematically derived from the bulk and edge spectral properties of the magnon-polaron bands that are characterized by their corresponding Chern numbers. Thereafter, we investigate transitions from one topological phase to another solely via tuning the spin-phonon coupling strength. Moreover, these transitions significantly impact the behavior of the thermal Hall conductivity that aids in discerning distinct topological phases. Additionally, the explicit dependencies on the temperature and the external magnetic field are explored in inducing topological phase transitions associated with the magnon-polaron bands. Thus, our work serves as an ideal platform to probe the interplay of frustrated magnetism and polaronic physics.
@@ I INTRODUCTION
Of late, geometrically frustrated kagome antiferromagnets (KAFM) have been proven to be an important candidate to host exotic QSL features [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. Characterized by various elementary excitations, the theoretical studies have speculated on ground state QSL markers of the kagome Heisenberg antiferromagnet, such as the $\mathbb{Z}_{2}$ [45, 36, 37, 41, 46], Dirac [39, 42, 43, 44], topologically entangled [38], and chiral [47, 24] spin liquid states, and more recently, spinon pair-density wave state [48]. Although, among the QSLs chiral spin liquids break the time-reversal symmetry (TRS) and result in formation of Chern bands, it is unclear whether the QSL ground state of the KAFM is gapped [35, 36, 37, 41, 37, 39, 43]. In addition to these fundamental issues, a KAFM harbors worthwhile quantum signatures induced by DMI [49, 50, 51, 52, 53, 54, 20, 21, 22, 23, 24]. The DMI can intrinsically induce long-range order (LRO), commonly termed $\bm{Q}=0$ magnetic order in frustrated kagome magnets, which is otherwise inaccessible at low temperatures. Hence, up to a critical order, DMI suppresses the QSL phase of the KAFM [51]. Past studies have evidenced the LRO in coplanar or noncollinear frustrated KAFM at specific regimes of temperature [55, 56, 49, 57, 50, 51]. However, theoretical predictions on non-coplanar KAFM structures, such as pyrochlore thin film [58], jarosites [59, 60, 61, 56, 49, 50], namely $\text{KFe}_{3}\text{(OH)}_{6}(\text{SO}_{4})_{2}$, $\text{AgFe}_{3}\text{(OH)}_{6}(\text{SO}_{4})_{2}$, etc., Vesignieite $\text{BaCu}_{3}\text{V}_{2}\text{O}_{8}\text{(OH)}_{2}$ [53], herbertsmithite $\text{ZnCu}_{3}\text{(OH)}_{6}\text{Cl}_{2}$ [62], and noncollinear canted configurations induced by a magnetic field [63, 64, 65] have been probed to discern MTHE experimentally. Due to the inherent noncoplanarity in several of these KAFMs, a DMI-induced scalar spin chirality, $\chi=\bm{S}_{i}.(\bm{S}_{j}\times\bm{S}_{k})$ originates which instills a nontrivial magnon topology [49, 50, 51, 54, 20, 21] and MTHE [20, 21, 22, 23, 24]. While a field-induced non-coplanar spin chirality is solely capable of manifesting MTHE in KAFMs, with no mandatory requirement of DMI [20]. In contrast, with no magnetic field, the in-plane DMI has been reported to stabilize a canted KAFM configuration with a finite spin chirality that significantly modifies the magnon thermal Hall conductivity [21] corresponding to each of the Chern (magnon) bands. Moreover, KAFMs host a “spinon quantum spin Hall effect” in the time-reversal symmetric gapped QSL phase and facilitate a continuous phase transition into this phase with the inclusion of an extended nearest neighbour hopping that melts the DMI-induced $\bm{Q}=0$ ordering [24]. Hence, this gapped QSL phase with oppositely moving spinon chiral edges (possessing spin Chern numbers, $C_{\uparrow,\downarrow}=\pm 1$) renders itself as a magnetic analogue of the $\mathbb{Z}_{2}$ topological insulator. However, even with all of these proposals being accounted for, there is still enough scope to explore quasiparticle excitonic topology and MTHE of the KAFM magnon bands.
@@
Intriguingly, other variants of quasiparticles, such as phonons (associated with the lattice vibrations), also corroborate nontrivial topological phenomena. Akin to phonon-induced band topology in electronic systems [66, 67, 68, 69, 70, 71, 72, 73], these vibrational modes can hybridize with magnons, resulting in magnon-phonon (also termed magnon-polaron) quasiparticle coupled topological signatures in magnetically ordered systems [74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93]. To mention a few, the hybridization effects of the magnon and phonon bands are mostly explored in square [77, 79, 82], honeycomb [81, 86, 87, 89, 90, 91] lattices, and two-dimensional (2D) antiferromagnets [75, 76, 79, 82, 86, 88, 90]. It has been claimed that hybridization of these quasiparticles can occur with (via breaking the mirror symmetry) [80, 77, 81, 86, 89] or without [78, 76, 74] DMI or through dipolar interactions [74], which results in anticrossing of these excitonic bands accompanied by a nontrivial Berry curvature that induces magnon-polaron thermal Hall effect. The hybridization in the absence of any DMI is usually explained by magnetostriction or magnetoelastic wave [94, 78, 76], which forms a magnon-polaron state mainly through the acoustic phonon modes. While the role of optical phonon modes has also been examined in manipulating the Chern numbers of the hybridized magnon-polaron bands [84, 86], such studies are non-existent in the context of a kagome lattice. Moreover, magnon-phonon topological physics in the kagome structure is not well established in the literature, especially on a KAFM configuration. Hence, a closer inspection of the magnon-optical phonon coupled quasiparticle scenario (in terms of magnon-polaron bound state) on a KAFM may unravel important topological properties of frustrated magnetism, which, to the best of our knowledge, has not been reported yet.
@@
In this study, we explore the effects of magnon-polaron excitations on a 2D non-coplanar KAFM (as schematized in Fig. 1(a)) in the presence of an out-of-plane DMI ($D$) and an external magnetic field ($B_{0}$). Additionally, we formulate our magnon model with the nearest neighbour (NN) and the next-nearest neighbour (NNN) Heisenberg exchange terms, namely $J_{1}$ and $J_{2}$, respectively. However, we mention that the in-plane DMI components ($D_{x,y}$) being negligibly small due to the dominant out-of-plane DMI ($D_{z}$) in most kagome materials [55, 62, 56, 51] we drop the in-plane DMI terms here. To inspect a frustrated magnetism on our system, the noncoplanarity or the canted configuration (away from $120^{\circ}$ N’eel ordering) of KAFM necessitates an externally applied out-of-plane magnetic field that solely generates a nontrivial magnon topology and MTHE [20]. In general, the interaction between quasiparticles in such a KAFM model can be intricate and hence requires involved techniques to derive an effective magnon picture. Here, we wish to investigate how the onsite (local) magnon modes couple to the optical phonon modes, resulting in a magnon-polaron state, and examine its impact on the magnon topology and the relative thermal Hall features of the entire KAFM structure. Focusing on this primary goal, we first consider that the optical phonon modes can couple with the spin (magnon) modes through two distinct mechanisms: (i) magnon modes at a particular lattice site interact with the same onsite phonon modes [95], named “local spin-phonon” coupling with a strength $g^{l}$, and (ii) local magnon interaction that involves phonons present at two NN lattice sites [95, 84, 78], dubbed “non-local spin-phonon” coupling with a strength $g^{nl}$. The purpose of dealing with optical (rather than the acoustic) phonon branches is to incorporate a stronger perturbation on the local atomic environment, and hence it ascertains more efficient coupling between the spin excitations and phonons [84], which can be probed via photo-excitation experiments. Furthermore, going beyond the adiabatic (high frequency) regime enables phonons to acquire a finite energy that stabilizes the KAFM configuration even at zero temperature, although a constraint on the phonon frequency should be imposed to avoid instability of the magnon-polaron bands. In addition, as optical modes are largely non-dispersive, it is convenient to model the phonon bands generically at small momenta [96] that are relevant for gap closing transitions of the KAFM bands. These justify the rationale for taking momentum-independent optical phonons into account. However, analytical methods for tackling such a coupling in a KAFM have not been proposed yet. Therefore, here we provide an analytical approach in the spirit of the spin-Peierls transformation method [95] to yield a finite temperature magnon-polaron hybridized bound state by completely decoupling these two distinct couplings on a KAFM, and then analyze numerically its topological properties and the thermal Hall conductivity. Although hitherto proposals on the kagome ferromagnet [81] and the kagome spin ice [93] have explained the effects of phononic excitations, the salient topology of the magnon-phonon hybridization in a frustrated geometry, such as a KAFM, has not been addressed till now. Therefore, the present study may hint at a richer topological scenario and bridge the gap between frustrated magnetism and magnon-polaron quasiparticle physics. Thus far, the following pertinent questions arise: (i) Can a topological transition be induced on the KAFM magnon-polaron bands solely by tuning the spin-phonon coupling strength? (ii) How will the magnon-polaron band topology differ in local and non-local interaction regimes, both qualitatively and quantitatively? (iii) What will the response of the topological thermal Hall conductivity be to such magnon-polaron hybridization in both these regimes?
@@
To address these points, we organize the remainder of the paper as follows. We begin by mathematically formulating the non-coplanar KAFM magnon model, and construct the momentum ($k$) space magnon-polaron Hamiltonian, followed by the exact decoupling of the aforementioned coupling through a canonical transformation in Sec. II. Subsequently, in Sec. III, we describe the numerical results in the following sequence. The bulk spectral properties accompanied by relative transitions in the Chern number are examined in Sec. III.1 via varying both the local and non-local coupling strengths distinctly, while the edge state features of the topological magnon-polaron bands are analyzed in Sec. III.2. Thereafter, the nature of the topological thermal Hall conductivities exhibiting the transitions is elaborately discussed in Sec. III.3. Finally, we summarize and conclude our findings in Sec. IV.
@@
Figure 1: Schematic diagram of a kagome antiferromagnet is displayed in (a), where red, blue, and green arrows refer to the spin orientation of A, B, and C sublattices, respectively, and $\mathbf{\delta}_{1}$, $\mathbf{\delta}_{2}$ and $\mathbf{\delta}_{3}$ denote three lattice vectors. (b) The direction of the out-of-plane DMI is shown, which acts downward from site 1 to 2, and upward from site 2 to 3, while (c) illustrates that in the presence of an external magnetic field, the spins become canted with an angle $\eta$ with respect to the $x$-$y$ plane.
@@ II.1 Magnon Hamiltonian
where $J_{1}(>0)$ and $J_{2}(>0)$ are the NN and the NNN antiferromagnetic exchange interaction strengths, as mentioned earlier. Additionally, $D$ is the coefficient of an out-of-plane DMI acting between two neighbouring spins, where $\nu_{ij}=\mp 1$ depending on their lattice positions, that is, negative (positive) for an upper (lower) triangle (see Fig. 1(b)) [20]. Finally, the Zeeman term involves the external transverse magnetic field $B_{0}$ that eventually generates the canted orientations of spins and further aids in inducing the nontrivial topology in the system. The lattice vectors connecting two same sublattice sites are denoted by, $\delta_{1}=a_{0}(1,\sqrt{3})$, $\delta_{2}=a_{0}(1,-\sqrt{3})$, $\delta_{3}=a_{0}(2,0)$.
Further, we consider non-collinear spin configurations with antiferromagnetic environment, where, neglecting all the spin fluctuations, the classical spin vectors at three different sublattice sites (A, B, C, as displayed in Fig. 1(a)) are given by
@@
Here, $\eta$ denotes the canting angle between the spin vectors and the lattice plane, as shown in Fig. 1(c). In the absence of an external magnetic field, the spin configuration exhibits a $120^{\circ}$ Néel (coplanar) order, with $\eta=0$. However, in the presence of an external magnetic field ($B_{0}\neq 0$), $\eta$ acquires a finite value, which can be determined by minimizing the classical energy ($\partial E_{\text{cl}}/\partial\eta=0$) corresponding to $H_{\text{s}}$ (Eq. (2)). This yields the relation $\sin{\eta}=B_{0}/B_{s}$, with $B_{s}$ being the saturation magnetization given by, $B_{s}=2\sqrt{3}[D+\sqrt{3}(J_{1}+J_{2})]$, corresponding to the ground-state spin configuration as reported in literature [20, 21].
Due to this canted spin orientation, the spin quantization axes at different sublattice sites not only differ from one another but also deviate from the global $z$-axis. Therefore, to establish a relation between the global spin operator $\mathbf{S}_{i}$ and the local spin operator $\tilde{\mathbf{S}}_{i}$, we need to introduce a sublattice-dependent rotation matrix, which can be expressed as
@@
Here $\tilde{S}_{i}^{+}=\tilde{S}_{x}+i\tilde{S}_{y}$, $\tilde{S}_{i}^{-}=\tilde{S}_{x}-i\tilde{S}_{y}$ and $a_{i}^{\dagger}(a_{i})$ is the creation (annihilation) operator corresponding to the A sublattice sites. Similarly, for the B and C sublattice sites, the creation (annihilation) operators are denoted by $b_{i}^{\dagger}$ ($b_{i}$) and $c_{i}^{\dagger}$ ($c_{i}$), respectively. In our case, these operators follow the same transformation as given in Eq. (5). In the frustrated antiferromagnetic system, the spin orientations at the different sublattice sites are specified in Eq. (4). By applying the spin-boson transformation defined in Eq. (5) to the spin Hamiltonian in Eq. (2), the Hamiltonian $H_{\text{s}}$ can be expressed in the magnon basis, and is given by
$H_{\text{s}}$ $=$ $S\sum_{r}\biggl{[}4(J_{1}+J_{2})\mathcal{E}+4D{\mathcal{E}}_{1}+\mu_{0}B_{0}\sin{\eta}\biggr{]}$ (6)
$\left(a_{r}^{\dagger}a_{r}+b_{r}^{\dagger}b_{r}+c_{r}^{\dagger}c_{r}\right)+\biggl{[}J_{1}\!(P\!+\!iQ)\!-\!D(P_{1}\!+\!iQ_{1})\biggr{]}$
$(c_{r}^{\dagger}a_{r}+c_{r}^{\dagger}a_{r+\delta_{1}}+b_{r}^{\dagger}c_{r}+b_{r}^{\dagger}c_{r+\delta_{2}}+a_{r}^{\dagger}b_{r}+a_{r+\delta_{3}}^{\dagger}b_{r})$
$+\biggl{[}J_{2}(P+iQ)\biggr{]}(a_{r}^{\dagger}b_{r-\delta_{2}}+a_{r+\delta_{1}}^{\dagger}b_{r}+b_{r+\delta_{1}}^{\dagger}c_{r}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+b_{r}^{\dagger}c_{r+\delta_{3}}+c_{r+\delta_{2}}^{\dagger}a_{r}+c_{r}^{\dagger}a_{r+\delta_{3}})$
$+\biggl{[}J_{1}R-DR_{1}\biggr{]}(c_{r}^{\dagger}a_{r}^{\dagger}+c_{r}^{\dagger}a_{r+\delta_{1}}^{\dagger}+b_{r}^{\dagger}c_{r}^{\dagger}+b_{r}^{\dagger}c_{r+\delta_{2}}^{\dagger}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+a_{r}^{\dagger}b_{r}^{\dagger}+a_{r+\delta_{3}}^{\dagger}b_{r}^{\dagger})$
$+J_{2}R(a_{r}^{\dagger}b_{r-\delta_{2}}^{\dagger}+a_{r+\delta_{1}}^{\dagger}b_{r}^{\dagger}+b_{r+\delta_{1}}^{\dagger}c_{r}^{\dagger}+b_{r}^{\dagger}c_{r+\delta_{3}}^{\dagger}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+c_{r+\delta_{2}}^{\dagger}a_{r}^{\dagger}+c_{r}^{\dagger}a_{r+\delta_{3}}^{\dagger})+h.c.,$
where $r$ denotes the position of a unit cell and the sum runs over all the unit cells. The operators $a_{r}^{\dagger}$ $(a_{r})$, $b_{r}^{\dagger}$ $(b_{r})$ and $c_{r}^{\dagger}$ $(c_{r})$ are the creation (annihilation) operators corresponding to sublattices A, B, and C, respectively, located at position $r$, corresponding to the triangle highlighted in yellow in Fig. 1(a). Moreover, each unit cell contains six bonds, comprising one upper and one lower triangle associated with that unit cell. Further, the remaining quantities of Eq. (6) are defined as